Find Q if the rate is 6% per year compounded yearly if outflows and inflows for...

  1. Find Q if the rate is 6% per year compounded yearly if outflows and inflows for a project over 5 years starting with Year 0 are as follows:
    1. Outflows: -$Q; -$18Q; -$25Q; $0; $0; $0
    2. Inflows: $0; $0; $0; $25,000; $85,000; $125,000
Answer #1

Assuming that the NPV will become 0 at 6% discount rate we can calculate the value of Q, for this we can use the goal seek function on excel:

Using it we get Q = 4517.05 and following schedule shows that NPV = 0

Year CF Discount Factor Discounted CF
0 $        -4,517.05 1/(1+0.06)^0= 1 1*-4517.0520365773= $         -4,517.05
1 $      -81,306.94 1/(1+0.06)^1= 0.943396226 0.943396226415094*-81306.9366583913= $       -76,704.66
2 $ -1,12,926.30 1/(1+0.06)^2= 0.88999644 0.88999644001424*-112926.300914432= $   -1,00,504.01
3 $        25,000.00 1/(1+0.06)^3= 0.839619283 0.839619283032302*25000= $        20,990.48
4 $        85,000.00 1/(1+0.06)^4= 0.792093663 0.79209366323802*85000= $        67,327.96
5 $    1,25,000.00 1/(1+0.06)^5= 0.747258173 0.747258172866057*125000= $        93,407.27
NPV = Sum of all Discounted CF $                        -